S/N 2 In Line PDBS w/o BS & Collet on output fiber

REVISION DATE: 03-Apr-2018 16:49:29

CONFIG 12: 23-Mar-2018 to 28-Mar-2018

In Line PSBS custom made by Mark, aligned on laser bench Pier 35, S/N 2. Test of temperature/humidity stability of Collet Fiber Connector. The collet output fiber labeled ORANGE is on the output of the through PDBS; input fiber still FISH #31. We took out the PDBS and the fold mirror. PD2 is connected to the through channel and PD1 has a 9V battery. FISH Fibers, SIRCUS Sphere, Yuqin Zong's white LED source, monitor with Si PD and SEI.

* E1 = Day 1: 23-Mar-2018 22degC, meant for 22/30/22 at 30%RH but forgot to turn on humidity control so this is just 22deg C.

* E2 = Day 2: 26-Mar-2018 22/30/22degC, 30% RH; also took darks afterwards.

* E3 = Day 3: 27-Mar-2018 22/30/22degC, 30% RH; darks before and after; FLUKE chl 2 moved close to Chamber temp/RH sensors.

* E4 = Day 4: 28-Mar-2018 22degC, 30%/40%/30% RH; FLUKE chl 2 moved back to near PDBS.


image

Figure 1: All data in this Configuration to date. PD signals are normalized to the monitor. Ratio PD1/Mon, Monitor, and normalized Signals all normalized to the initial value.


image

Figure 2: Zoomed in look at Day 1


image

Figure 3: Zoomed in look at Day 2


image

Figure 4: Zoomed in look at Day 3


image

Figure 5: Zoomed in look at Day 4



pwd: \\elwood.nist.gov\68_PML\685\internal\Restricted\Johnson\MOBY-AHAB\fiber splitter tests\Config12
Date: 03-Apr-2018 16:49:29
Created from plt_output_panels_Config12_(1)